OpenStudy (aliaaalihassan):

complete the square and explain how to transform the graph of y = x^2 into the graph of the given function f(x)= x^2 +2x+4

OpenStudy (cwrw238):

x^2 + 2x + 4 = (x + 1)^2 - 1 + 4 = ( x + 1)^2 + 3

OpenStudy (cwrw238):

the + 2 in the brackets transforms x^2 by moving the while curve 2 units to the left parallel to the x-axis the +3 moves the graph 3 units upwards parallel to y-axis

OpenStudy (cwrw238):